# Fair Ordering via Streaming Social Choice Theory

### v3

@@ 5. Ranked Pairs Voting
We turn now to the Ranked Pairs method (tideman1987independence). Ties in edge weights are broken according to an arbitrary (deterministic) rule.
Input: An ordering vote from each replica $\{\sigma_{i}\}_{i\in[n]}$
$G=(V,E,w)\leftarrow G(\sigma_{1},...,\sigma_{n})$ ;
  /* Compute ordering graph */
$H\leftarrow(V,\emptyset)$ ;
  /* Initialize empty graph on same vertex set */
Sort $E$ by edge weight
foreach $e\in E$ do
       Add $e$ to $H$ if it would not create a directed cycle in $H$
end foreach
Output: The topological sort of $H$
  /* The loop considered every edge, so the topological sort of $H$ is unique */
ALGORITHM 1 Ranked Pairs Voting
@@
Lemma 4.2 and Theorem 5 together imply an important property. The output of Ranked Pairs can be divided into batches consistent with $\gamma$-batch-order-fairness for any $\gamma$, not just a single $\gamma$ (subject to the interleaving discussed in §4.2). Unlike prior work, there is no need to choose a $\gamma$ or order transactions within a batch.
@@ 6. Streaming Ranked Pairs
Our algorithms must be monotonic (Definition 6.1)—if replicas extend their ordering votes, the algorithm, when run again on the extensions, can only extend its prior output.
@@
Observe that if a sequencing algorithm is monotonic, it implies a well-formed definition for aggregating a set of orderings on countably infinite sets of transactions, not just on finite inputs. Specifically, $\tau$ comes before $\tau^{\prime}$ in the overall (infinite) case if there exists a finite subset of the input orderings such that the algorithm puts $\tau\prec\tau^{\prime}$ in its output.
@@ 6.1. Decisions Are Local
If the path visits a transaction $\tau$ in $R_{\tau_{A}}$, then it creates a cycle with the edge from $\tau_{A}$ to $\tau$ (which ranked pairs must choose, by Observation 6.1). If the path visits a transaction $\tau$ in $P_{\tau_{B}}$, then it creates a cycle with the edge from $\tau$ to $\tau_{B}$. ∎
@@
First, we construct an ordering graph that captures the information available to an algorithm at a finite time.
@@
Suppose that each replica submits an ordering vote $(\sigma_{1},...,\sigma_{n})$. Let $V$ be the set of all transactions that appear in each vote, and let $\hat{v}$ be the “future” vertex.
The Streamed Ordering Graph is a weighted, directed, complete graph $\hat{G}=(V^{\prime},E)$ on vertex set $V\cup\{\hat{v}\}$.
@@ 6.2. A Streaming Algorithm
We can now construct an algorithm to compute Ranked Pairs in our streaming setting. Instead of either including or rejecting every edge, Algorithm 2 has the option to declare an edge as indeterminate; that is, the algorithm is not able to determine whether to include or exclude the edge with the information in its input.
Input: An ordering vote from each replica $\{\sigma_{i}\}_{i\in[n]}$
$\hat{G}=(V,E,w)\leftarrow\hat{G}(\sigma_{1},...,\sigma_{n})$ ;
  /* Compute streamed ordering graph */
$H\leftarrow(V,\emptyset)$ ;
  /* Initialize empty graph on same vertex set */
foreach edge $(\tau,\hat{v})$ and $(\hat{v},\tau)$ with $w(e)=1$ do
       Add $e$ to $H$, and mark it as indeterminate
end foreach
Sort $E$ by edge weight, with a fixed tiebreaking rule.
foreach $e=(\tau_{i},\tau_{j})\in E$ do
        /* Recall that $\hat{v}$ is the ‘‘future’’ vertex of the streamed ordering graph */
       If there is a directed path in $H$ restricted to $U_{\tau_{i},\tau_{j}}$ from $\tau_{j}$ to $\tau_{i}$ where every edge is determinate, then do not include the edge in $H$.
       If there is no directed path in $H$ restricted to $U_{\tau_{i},\tau_{j}}$ from $\tau_{j}$ to $\tau_{i}$ where every edge is either determinate or indeterminate, add $(\tau_{i},\tau_{j})$ to $H$ and mark it as determinate.
       Otherwise, add $(\tau_{i},\tau_{j})$ to $H$ and mark it as indeterminate.
end foreach
Output: The topological sort of $H$ (with sorting comparisons restricted to determinate edges), up to but not including the first transaction $\tau$ bordering an indeterminate edge
ALGORITHM 2 Streaming Ranked Pairs
To prove correctness of this algorithm, it suffices to show that the output of the algorithm is monotonic and that the output is consistent with the non-streaming version of Ranked Pairs. By consistency, we specifically mean that at any point, if clients were to stop sending new transactions and all replicas eventually voted on every replica, Algorithms 1 and 2 would produce the same output.
@@ 7. Liveness and Efficiency
ALGORITHM 3 Streaming Ranked Pairs, Version 2
@@ 7.2. Protocol Liveness
Algorithm 2 is not asymptotically live. §14 gives an instance on which the algorithm would never output any transaction.
That construction uses an adversarial tiebreaking rule, but Ranked Pairs requires only some tiebreaking rule. Our algorithm can therefore generate a tiebreaking rule dynamically. Specifically, when the algorithm visits edges of weight $\gamma=\frac{k}{n}$, it defers edges that would become indeterminate to the end of the ordering among edges of that weight.
Algorithm 3 constructs a (partial) ordering between edges that breaks ties between edges of the same weight. The output of this algorithm is, by the same arguments as in §6.2, consistent with the output of non-streaming Ranked Pairs, when using an tiebreaking rule consistent with this edge ordering. The oracle mechanism of §7.1 passes this tiebreaking information from one run of Algorithm 3 to the next.
@@
Crucially, this small change enables the following structural lemma. We start with a useful definition.
Definition 7.1 ().
@@
By contrast, the best bound in Algorithm 2 is that there exists a contemporary indeterminate edge of weight $\frac{k}{n}$.
@@
Lemma 7.2 implies that for any indeterminate edge, there is sequence of contemporary indeterminate edges of strictly increasing weights that causes the edge to be indeterminate.
@@
Additionally, under network synchrony, Lemma 7.2 bounds the temporal delay between when a transaction is sent by a client and when Algorithm 3 adds it to its output.
Assumption 7.2 (Synchronous Network).
@@
As such, replicas can disagree on the ordering between $\tau$ and $\tau^{\prime}$ only if they were sent at roughly the same time.
@@
By construction, $\tau\notin R_{\tau_{i}}$, so $t\leq t_{i}+\Delta$. Furthermore, $\tau\notin P_{\tau_{j}}$, so $t+\Delta\geq t_{j}$. Furthermore, by Observation 7.2, $t_{j}+\Delta\geq t_{i}$.
@@ 7.3. Trading Accuracy for Liveness
8. Protocol Instantiation
The algorithms thus far have assumed that every replica will eventually vote on every transaction. Any faulty replica, however, could exclude a transaction from its output, leaving the transaction (and the protocol) permanently stuck.
The main idea to resolve this difficulty is that if a replica refuses to vote on a transaction, then that replica must be faulty. If a replica is faulty, then it suffices for the rest of the replicas to make up a vote on its behalf.
We give here an example method for instantiating our sequencing protocol on top of existing protocol components, although this is far from the only method.
As a base layer, we require some consensus protocol, such as (yin:hotstuff; lokhava:stellar; castro1999practical; miller2016honey), that proceeds in rounds. At any time, every replica can submit a continuation of its previous ordering; that is, an ordered list of transactions that extends its ordering vote. The consensus protocol then comes to consensus on a set of received ordering votes. The Streaming Ranked Vote algorithm is run (by every replica) on the set of received votes after every new set of votes is finalized by the consensus protocol. Assuming correctness of the consensus protocol, every replica will run the algorithm on the same input, and thereby produce the same output ordering.
We assume that every replica has a registered public key. Every vote continuation must be cryptographically signed by the associated secret key. Furthermore, every continuation must be accompanied by the hash of the previously submitted continuation, so that if one continuation vote from a replica is lost (perhaps by malicious activity of consensus participants) the replica’s ordering cannot be manipulated.
We make a network synchrony assumption, akin to Assumption 7.2. If a transaction appears first in some replica’s vote in round $T$ of the consensus protocol, then we assume that there is some known bound $k$ such that every other honest replica will have time to receive and submit a vote on the transaction before the end of round $T+k$. As such, after every round $r$, the protocol computes the set of all transactions that have not received votes from every replica but were voted on by some replica at or before round $r-k$. For each of these transactions $\tau$, the protocol appends a vote for $\tau$ at the end of a replica’s current vote sequence if $\tau$ is not already present. When appending multiple transactions to a replica’s vote, transactions can be sorted deterministically (i.e. by hash).
This protocol guarantees that every transaction will receive a vote from all replicas within at most $k$ rounds after it is first submitted to the protocol. In this setting, $k$ must be configured in advance by the protocol. The parameter must be set so that every honest replica is able to submit a vote on a transaction within $k$ rounds. Naturally, the choice of $k$ depends on network conditions the censorship-resistance (or “chain-quality” (garay2015bitcoin)) of the consensus protocol.
As an example alternative, one could also use the protocol in kelkar2020order, which is based on atomic broadcast and a “Set Byzantine Agreement” primitive. This protocol waits until all honest replicas are able to add a transaction $\tau$ to their ordering votes, and then invokes a protocol to agree on the set of replicas $U_{\tau}$ which have the transaction in their ordering votes. This set may not be the set of all replicas, but includes all honest replicas. To use our ordering protocol on this consensus protocol, it suffices to make up a vote for the (presumed faulty) replicas not in $U_{\tau}\cap U_{\tau^{\prime}}$. Replicas in $U_{\tau}\setminus U_{\tau^{\prime}}$ can be presumed to vote $\tau\prec\tau^{\prime}$ and vice versa, and replicas not in $U_{\tau}\cup U_{\tau^{\prime}}$ can be presumed to vote (arbitrarily) by, e.g., comparing deterministic hashes of $\tau$ and $\tau^{\prime}$.

### v4

@@ 4. Ranked Pairs Voting
Our protocol is based on the Ranked Pairs method (tideman1987independence). Ranked Pairs (Algorithm 2, reproduced in §15) first builds the ordering graph (Definition 3.2), then iterates through edges in order of edge weight (breaking ties arbitrarily). It builds an acyclic set of edges greedily, adding each new edge so long as it does not create a cycle with edges already in the set. The output ordering is the result of topologically sorting the resulting acyclic graph.
@@
Lemma 3.2 and Theorem 4 together imply an important property. Ranked Pair’s output can be divided into batches consistent with $\gamma$-batch-order-fairness for any $\gamma$ (subject to the interleaving discussed in §3.2). There is no need to choose a $\gamma$ or arbitrarily order transactions within a batch.
@@ 5. Streaming Ranked Pairs
An algorithm must be monotonic (Definition 5.1)—if replicas extend their ordering votes, the algorithm, when run again on the extensions, can only extend its prior output.
@@
If a sequencing algorithm is monotonic, it implies a well-formed definition for aggregating a set of orderings on countably infinite sets of transactions, not just on finite sets. $\tau$ comes before $\tau^{\prime}$ in the infinite case if there exists a finite subset of the input orderings such that the algorithm puts $\tau\prec\tau^{\prime}$ in its output.
@@ 5.1. Decisions Are Local
If the path visits a transaction $\tau$ in $R_{\tau_{A}}$, then it creates a cycle with the edge from $\tau_{A}$ to $\tau$ (which ranked pairs must choose, by Observation 5.1). If the path visits a transaction $\tau$ in $P_{\tau_{B}}$, then it creates a cycle with the edge from $\tau$ to $\tau_{B}$. ∎
@@
First, we construct an ordering graph that captures the information available at a finite time.
@@
Suppose that each replica submits an ordering vote $(\sigma_{1},...,\sigma_{n})$. Let $V$ be the set of all transactions that appear in each vote, and let $\hat{v}$ be a new vertex (the “future”) The Streamed Ordering Graph is a weighted, directed, complete graph $\hat{G}=(V^{\prime},E)$ on vertex set $V\cup\{\hat{v}\}$.
@@ 5.2. A Streaming Algorithm
We can now construct a streaming algorithm to compute Ranked Pairs. Instead of either including or rejecting every edge, Algorithm 3 (shown in full in Appendix 15) has the option to declare an edge as indeterminate; that is, the algorithm cannot determine whether to include or exclude the edge with the information in the input.
The algorithm proceeds identically to the classic Ranked Pairs algorithm, except regarding how it treats indeterminate edges. For each new edge, it first checks whether it can accept the edge (without creating a cycle of accepted edges), supposing that all currently indeterminate edges are later accepted. Conversely, it also checks whether it can reject the edge, which means checking whether accepting the edge would always create a cycle of accepted edges (assuming the currently indeterminate edges are later rejected). The algorithm operates on the modified input graph (Definition 5.4), where the “future” vertex bounds the information the algorithm has to consider about as-yet-unseen transactions. Informally, all edges between as-yet-unseen transactions are indeterminate, so from the perspective of directed connectivity analysis, the unseen transactions form a single, strongly-connected component.
To prove correctness of this algorithm, it suffices to show that the output of the algorithm is monotonic and that the output is consistent with the non-streaming version of Ranked Pairs. By consistency, we specifically mean that at any point, if clients were to stop sending new transactions and all replicas eventually voted on every replica, Algorithms 2 and 3 would produce the same output.
@@ 6. Liveness and Efficiency
ALGORITHM 1 Streaming Ranked Pairs
@@ 6.2. Protocol Liveness
Algorithm 3 is not asymptotically live. Appendix 13 gives a worst-case example. The problem is the tiebreaking rule for how Ranked Pairs iterates over edges of equal weight. That construction uses a worst-case tiebreaking rule, but Ranked Pairs requires only some tiebreaking rule. Our algorithm can generate a tiebreaking rule dynamically. Specifically, when the algorithm visits edges of weight $\gamma=\frac{k}{n}$, it defers edges that would become indeterminate to the end of the ordering among edges of that weight.
Algorithm 1 thus constructs a (partial) ordering between edges that breaks ties between edges of the same weight. The output of this algorithm is, by the same arguments as in §5.2, consistent with the output of non-streaming Ranked Pairs, when using an tiebreaking rule consistent with this edge ordering. The oracle mechanism of §6.1 passes this tiebreaking information from one run of Algorithm 1 to the next.
@@
This small change enables the following structural lemma. We start with a useful definition.
Definition 6.1.
@@
By contrast, the best bound in Algorithm 3 is that there exists a contemporary indeterminate edge of weight $\frac{k}{n}$. Lemma 6.2 implies that for any indeterminate edge, there is sequence of contemporary indeterminate edges of strictly increasing weights that causes the edge to be indeterminate.
@@
Additionally, given a network assumption like Assumption 6.2, Lemma 6.2 bounds the temporal delay between when a transaction is sent by a client and when Algorithm 1 adds it to its output.
Assumption 6.2 (Network Delay Bound).
@@
As such, replicas disagree on the ordering between $\tau$ and $\tau^{\prime}$ only if they were sent at roughly the same time. Note that Algorithm 1 does not require knowledge of $\Delta$, and does not depend on Assumption 6.2 for correctness.
@@
By construction, $\tau\notin R_{\tau_{i}}$, so $t\leq t_{i}+\Delta$. Furthermore, $\tau\notin P_{\tau_{j}}$, so $t+\Delta\geq t_{j}$. Furthermore, by Observation 6.2, $t_{j}+\Delta\geq t_{i}$.
@@ 6.3. Trading Accuracy for Liveness
7. Related Work
7.1. Order-Fairness
Closest to this work are Aequitas (kelkar2020order) and Themis (themis), which propose the definition of $\gamma$-batch-order-fairness and instantiate ordering protocols based on aggregating ordering votes into a total order. Aequitas achieves $(\gamma,\frac{f}{n})$-minimal-batch-order-fairness for a fixed choice of $\gamma$, but is not asymptotically live. aequitaspermissionless instantiate Aequitas in a setting where the number of replicas is not known, and similarly points out the connection between this streaming problem and the classical ranking aggregation problem. §3.3 gives a more detailed comparison to these works.
cachin2022quick give a notion of “$\kappa$-differential-order-fairness;” Theorem D.1 of (themis) shows equivalence of this definition and $\gamma$-batch-order-fairness, for a choice of $\gamma$ (as a function of $\kappa$).
Our definition of $(\gamma,\delta)$-minimal-batch-order-fairness (Definition 3.4) strengthens the (vacuously satisfiable) $\gamma$-batch-order-fairness definition used in prior work (see §3 for details).
Also concurrently with this work, kiayias2024ordering propose a different method for ordering transactions within batches (but rely on a specific choice of a parameter $\gamma$ to determine those batches). A version of that protocol adds timestamps on reported orderings to improve liveness.
Additionally, Themis’s (themis) round-based sequencing protocol presents a different problem insurmountable by any process like the one suggested by vafadar2023condorcet that sorts batches separately. Namely, Themis may finalize the composition of a batch before has enough information to guarantee that, according to the Ranked Pairs ordering, all transactions in the batch must come before all those not in the batch (and not already in the output). Thus, it may output two batches in sequence, such that running Ranked Pairs on each batch separately produces a different output than running Ranked Pairs on the union of the two batches together. We give a example of this phenomenon in §12.
zhang2020byzantine add timestamps to ordering votes and sorts transactions by median timestamp to provide a linearizability guarantee, which is incomparable to $(\gamma,\delta)$-minimal-batch-order-fairness. mamageishvili2023buying propose letting users pay to reduce their reported timestamps.
7.2. Social Choice
7.3. Front-Running
The study of order-fairness is often motivated by the goal of limiting an adversary’s ability to order transactions (“front-running”) in public blockchains (as in, e.g., daian2020flash). Additional approaches to this problem include commiting to an ordering before any replica can know transaction contents, using threshold encryption or commit-reveal schemes (malkhi2022maximal; clineclockwork; zhang2022flash). Li et. al (li2023transaction) study the problem of verifiably computing an ordering within a trusted hardware enclave. These could be applied on top of a streaming ordering algorithm.
Kavousi et. al (kavousi2023blindperm) randomly shuffle blocks of encrypted transactions. Some decentralized exchanges (ramseyer2023speedex; cowswapproblem; penumbraswap) process transactions in unordered batches, eliminating the need for an ordering within a block (but not eliminating all order manipulations (zhangcomputation)). ferreira2022credible and li2023mev construct sequencing rules specific to decentralized exchanges.